# Differentiation Through Black-Box Quadratic Programming Solvers

### v3

@@ Abstract
Differentiable optimization has attracted significant research interest, particularly for quadratic programming (QP). Existing approaches for differentiating the solution of a QP with respect to its defining parameters often rely on specific integrated solvers. This integration limits their applicability, including their use in neural network architectures and bi-level optimization tasks, restricting users to a narrow selection of solver choices. To address this limitation, we introduce dQP, a modular and solver-agnostic framework for plug-and-play differentiation of virtually any QP solver. Our key theoretical insight is that the solution and its derivative can each be expressed in terms of closely-related and simple linear systems by using the active set at the solution. This insight enables efficient decoupling of the QP’s solution, obtained by any solver, from its differentiation. Our open-source, minimal-overhead implementation will be made publicly available and seamlessly integrates with more than 15 state-of-the-art solvers. Comprehensive benchmark experiments demonstrate dQP’s robustness and scalability, particularly highlighting its advantages in large-scale sparse problems.
@@ 1 Introduction
Computational methods often rely on solving optimization problems, i.e., finding an optimum of an objective function subject to constraints. This has led to the development of numerous high-performance open-source and commercial solvers, particularly for constrained convex optimization [Wright, 2006; Boyd & Vandenberghe, 2004]. Recently, there has been growing interest in differentiable optimization, which enables gradient-based learning through optimization layers that show promise across applications such as image classification [Amos et al., 2017], optimal transport [Rezende & Racanière, 2021; Richter-Powell et al., 2021], zero-sum games [Ling et al., 2018], tessellation [Chen et al., 2022], control [Amos et al., 2018; de Avila Belbute-Peres et al., 2018; Ding et al., 2024], decision-making [Tan et al., 2020], robotics [Holmes et al., 2024], biology [Zhang et al., 2023], and NLP [Thayaparan et al., 2022].
Differentiable optimization focuses on computing the derivatives of the optimal solution to an optimization problem with respect to parameters defining the problem’s objective and constraints. Rooted in classical sensitivity analysis and parametric programming Rockafellar [1970]; Rockafellar & Wets [1998]; Fiacco [1983, 1990]; Fiacco & McCormick [1968]; Lee et al. [2010]; Bonnans & Shapiro [2013]; Pistikopoulos et al. [2020], it has gained momentum through the idea that optimization problems can be embedded as parameterized “layers” within neural networks – training such layers requires computing the derivative (gradient) of the optimization problem’s solution during backpropagation [Amos & Kolter, 2017a; Agrawal et al., 2019a; Blondel & Roulet, 2024].
@@
Despite significant research efforts, existing approaches for differentiating QPs fail to fully leverage state-of-the-art solvers. General-purpose methods that support a broad class of optimization problems often include QPs, but they sacrifice efficiency and robustness for generality, underperforming compared to QP-specific methods. On the other hand, recent QP-specific approaches are tightly coupled with particular or proprietary solvers (e.g., for batching small QPs), limiting their flexibility [Amos & Kolter, 2017a; Bambade et al., 2024; Butler, 2023].
This tight integration restricts broader applicability. Solving QPs efficiently and reliably often requires advanced solvers like Gurobi [Gurobi Optimization, LLC, 2024] and MOSEK [Andersen & Andersen, 2000], developed over years of academic and commercial effort. These solvers handle scales and complexities that non-industrial implementations cannot. However, no single solver fits all problems, making solver choice critical. To our knowledge, no existing approach provides a robust, efficient, and fully flexible framework and implementation for differentiating QPs across solvers.
We introduce dQP, a modular, solver-agnostic framework that transforms any QP solver into a differentiable layer. Our key insight is that the gradient of a QP solution can be computed explicitly using the primal solution and the active set of constraints, both obtainable from the solver. Building on classical sensitivity theory, we show that the gradient can be expressed through a reduced linear system involving only the active constraints. Moreover, the active set can be used not only to simplify the gradient computation but also to easily recover the dual solution if not returned by the solver.
We implement dQP as an open-source, minimal-overhead layer on top of qpsolvers [Caron et al., 2024b], enabling seamless integration with over 15 commercial and free solvers, and straightforward extension to additional ones. Our modular approach allows users to select the best solver for their task while providing differentiability. We evaluate dQP on a benchmark of over 2,000 diverse QPs and demonstrate superior performance compared to existing methods, particularly in large-scale structured sparse problems, as demonstrated in Figure 1.
@@
    We prove that QPs can be explicitly differentiated using only the primal solution via a reduced-dimension locally equivalent linear system determined by the active set.
@@
    We design and implement a modular differentiable layer compatible with any QP solver, enabling plug-and-play flexibility. Our open-source code will be released publicly.
@@
Figure 1: Comparison of differentiable QP methods for projection onto the probability simplex, evaluated by (a) solve time and (b) solution accuracy (duality gap). For moderately sized problems, our approach, using the Gurobi QP solver, outperforms existing methods on both metrics. As problem size increases, our method remains efficient, while others become intractable.
@@ 2 Related Works
Differentiable Optimization. OptNet [Amos & Kolter, 2017a; Amos, 2019] differentiates QPs through their optimality conditions, and focuses on small dense problems for GPU batching. They solve the full Jacobian system efficiently by reusing the factorization employed in their custom interior-point method. However, as noted in [Bambade et al., 2024], this comes at the cost of ill-conditioning due to symmetrization. More recent differentiable QP methods include Alt-Diff and SCQPTH [Sun et al., 2022; Butler & Kwon, 2023; Butler, 2023] which use first-order ADMM and approximately differentiate the fixed point map, and QPLayer [Bambade et al., 2024] which accommodates infeasibility with extended conservative Jacobians. Similarly to OptNet, several of these are tightly integrated with custom algorithms, often to enable access to internal computations required for differentiation. Alt-Diff is coupled with a custom ADMM method, SCQPTH reimplements OSQP, and QPLayer is built on ProxQP. Several works have noticed the importance of the active constraint set in differentiating constrained optimization problems [Amos et al., 2017; Gould et al., 2022; Paulus et al., 2021], as well as in the context of quadratic programming [Amos et al., 2018; Bambade et al., 2024; Pan et al., 2024; Niculae et al., 2018]. A common observation is that the algebraic system obtained through implicit differentiation can be simplified by removing rows corresponding to inactive constraints. Some works have additionally observed that backpropagation itself can be cast as a distinct equality-constrained QP parameterized by the incoming gradients [Amos et al., 2017; Pan et al., 2024]. However, existing approaches have not used these observations to create a differentiable layer that supports arbitrary black-box QP solvers, missing the opportunity to decouple optimization and differentiation. Other classes of optimization problems, such as convex cone programs [Agrawal et al., 2019b] and mixed-integer programs [Paulus et al., 2021], have also been differentiated. Some frameworks [Agrawal et al., 2019a; Blondel et al., 2022; Pineda et al., 2022; DeepMind et al., 2020; Ren* et al., 2023; Besançon et al., 2024; Paulus et al., 2024; Schaller & Boyd, 2025] provide differentiable interfaces to broader classes of optimization problems, but their support of QP has significant limitations. CVXPYLayers [Agrawal et al., 2019a] reformulates the QP into a cone program to use diffcp internally [Agrawal et al., 2019b] and, as a result, does not support specialized QP solvers, relying exclusively on the cone solvers SCS, ECOS, and Clarabel. The framework Theseus [Pineda et al., 2022] directly handles only unconstrained problems and similarly lacks support for QP-specific solvers. TorchOpt [Ren* et al., 2023] and Optax [DeepMind et al., 2020] support unconstrained optimizers used in meta-learning, but require user-defined optimality conditions and user-supplied solvers to handle problems such as QPs. JAXopt [Blondel et al., 2022] includes an implicit differentiation wrapper for CVXPY, which requires symbolic compilation of the QP and does not support sparse matrices; for sparse QPs, JAXopt includes a differentiable re-implementation of OSQP (similar to SCQPTH). Altogether, these drawbacks lead to subpar performance which is reported for some generic methods in previous work on differentiable QPs (e.g., [Bambade et al., 2024]). For completeness, Table 4 in Appendix A summarizes 24 relevant methods, including those discussed above.
@@
Implicit Layers. Optimization layers are a class of implicit layers that use implicit differentiation to compute gradients of solution mappings that lack closed-form expressions [Duvenaud et al., 2020]. Another class of implicit layer are deep equilibrium models, which are defined by fixed-point mappings and can be interpreted as infinitely deep networks [Bai et al., 2019; Kawaguchi, 2021; El Ghaoui et al., 2021; Gurumurthy et al., 2021; Winston & Kolter, 2020; Bai et al., 2020]. Similar techniques extend beyond algebraic equations to neural ODEs, where the adjoint state method from parametric PDE control is applied [Lions, 1971; Xue et al., 2020; Beatson et al., 2020; Chen et al., 2018]. Implicit differentiation also plays a key role in bi-level programming [Colson et al., 2007; Kunisch & Pock, 2013; Gould et al., 2016; Alesiani, 2023] and meta-learning, where it enables optimization of the outer learning loop [Finn, 2018; Andrychowicz et al., 2016; Hochreiter et al., 2001; Hospedales et al., 2021; Rajeswaran et al., 2019; Sambharya et al., 2024]. Alternative methods bypass implicit differentiation using approximations – for example, by applying automatic differentiation to iterative algorithms via loop unrolling [Belanger & McCallum, 2016; Belanger et al., 2017; Metz et al., 2019; Scieur et al., 2022], or by differentiating a single iteration or using Jacobian-free techniques for fixed-point mappings [Geng et al., 2021; Fung et al., 2022; Bolte et al., 2023].
@@
Sensitivity Analysis and Parametric Programming. There is extensive mathematical theory on the local behavior of optimization problems under perturbations [Rockafellar, 1970; Rockafellar & Wets, 1998], particularly regarding solution sensitivity and stability [Fiacco, 1983, 1990; Fiacco & McCormick, 1968; Lee et al., 2010; Bonnans & Shapiro, 2013]. While the implicit function theorem is central to this analysis, applying it to optimization problems requires an intermediate step: reformulating the problem through its optimality conditions, which demands several regularity assumptions. Sensitivity theory supports applications such as multi-parametric programming [Pistikopoulos et al., 2020], including model predictive control, where problems must often be solved repeatedly for many parameter values, increasing computational costs. To address this, [Bemporad et al., 2002] observed that QPs admit closed-form solutions when the active set is known in advance, enabling offline precomputation. The parameter space can be partitioned into regions with fixed active sets [Spjøtvold et al., 2006], where the solution is stable under perturbations. This idea continues to inform modern methods [Ferreau et al., 2014; Narciso et al., 2022; Arnström & Axehill, 2024].
@@ 3 Approach
We begin by formulating the problem, establishing the basic theory of differentiation of QPs, connecting this to our main novel theoretical observation, and concluding with a straightforward algorithm. We note that various subparts of the background we discuss have been used to develop methods for differentiating QPs, see Section 2. However, no single work has fully developed the explicit theory we present or leveraged it to practically implement a fully modular differentiable QP layer.
@@ 3.1 Problem Setup: Differentiating QPs
This work focuses on computing $\partial_{\theta}z^{*}(\theta)=\frac{\partial}{\partial\theta}z^{*}(\theta)$, the derivative of the optimal point of the QP (1) with respect to the parameters $\theta$. Intuitively, this derivative quantifies the change in the optimal point of the QP in response to a perturbation of its parameters $\theta$. Our goal is to compute $\partial_{\theta}z^{*}(\theta)$ efficiently and independently of the method used to approximate the optimal point $z^{*}(\theta)$.
An important use case, highlighted in recent work on differentiable optimization, involves incorporating a QP of the form (1) as the $\ell$-th layer of a neural network. In this setting, the input to the QP layer, $x_{\ell}$, serves as the parameter vector $\theta$ that defines the QP’s objective and constraints, and the output is the optimal solution $x_{\ell+1}=z^{*}(x_{\ell})$. Generally, training such a network requires backpropagating gradients through each layer, which involves computing the Jacobian $\frac{\partial x_{\ell+1}}{\partial x_{\ell}}$. For a QP layer, this Jacobian corresponds to $\partial_{\theta}z^{*}(\theta)$, the derivative of the optimal point with respect to the problem parameters. This derivative is also crucial in descent-based methods for solving bi-level optimization problems Colson et al. [2007].
@@ 3.2 KKT Conditions and Sensitivity Analysis
The first-order Karush–Kuhn–Tucker (KKT) conditions [Karush, 1939; Kuhn & Tucker, 1951; Boyd & Vandenberghe, 2004; Wright, 2006] provide a useful algebraic characterization of the optimal points of constrained optimization problems. For the QP (1), the KKT conditions take the form,
@@
To differentiate QPs, we apply the Basic Sensitivity Theorem (Theorem 2.1 in [Fiacco, 1976]), which underpins differentiation of the KKT conditions with respect to $\theta$. Differentiability at $\theta$ requires the additional assumption of strict complementary slackness, ruling out the degenerate case where both $(Cz^{*}-d)_{j}=0$ and $\mu^{*}_{j}=0$, thus ensuring that the active set remains unchanged under small perturbations of $\theta$. Under this assumption, the primal-dual point $\zeta^{*}(\theta)=(z^{*}(\theta),\lambda^{*}(\theta),\mu^{*}(\theta))$ is differentiable in a neighborhood of $\theta$, optimal for the QP (1), uniquely satisfies the KKT conditions (2), and maintains strict complementary slackness. Crucially, the active set $J(\theta)$ remains fixed in this neighborhood.
@@
Under the assumptions of the Basic Sensitivity Theorem, the linear system (3) is invertible. It degenerates exactly in the presence of weakly active constraints $\mu_{j}^{*}=(Cz^{*}-d)_{j}=0$, for which the QP is non-differentiable (see, e.g., Amos & Kolter [2017a]). For any inactive constraint $j\notin J$, the dual variable $\mu_{j}^{*}$ vanishes, and thus the corresponding rows and columns of (3) can be removed, yielding the simplified reduced form,

### v4

@@ Abstract
Differentiable optimization has attracted significant research interest, particularly for quadratic programming (QP). Existing approaches for differentiating the solution of a QP with respect to its defining parameters often rely on specific integrated solvers. This integration limits their applicability, including their use in neural network architectures and bi-level optimization tasks, restricting users to a narrow selection of solver choices. To address this limitation, we introduce dQP, a modular and solver-agnostic framework for plug-and-play differentiation of virtually any QP solver. A key insight we leverage to achieve modularity is that, once the active set of inequality constraints is known, both the solution and its derivative can be expressed using simplified linear systems that share the same matrix. This formulation fully decouples the computation of the QP solution from its differentiation. Building on this result, we provide a minimal-overhead, open-source implementation (https://github.com/cwmagoon/dQP) that seamlessly integrates with over 15 state-of-the-art solvers. Comprehensive benchmark experiments demonstrate dQP’s robustness and scalability, particularly highlighting its advantages in large-scale sparse problems.
@@ 1 Introduction
Computational methods often rely on solving optimization problems, i.e., finding an optimum of an objective function subject to constraints. This has led to the development of numerous high-performance open-source and commercial solvers, particularly for constrained convex optimization [114, 27]. Recently, there has been growing interest in differentiable optimization, which enables gradient-based learning through optimization layers that show promise across applications such as image classification [8], optimal transport [97, 98], zero-sum games [75], tessellation [35], control [9, 38, 40], decision-making [110], robotics [62], biology [116], and NLP [111].
Differentiable optimization focuses on computing the derivatives of the optimal solution to an optimization problem with respect to parameters defining the problem’s objective and constraints. Rooted in classical sensitivity analysis and parametric programming [99, 100, 47, 49, 48, 73, 26, 93], differentiable optimization has gained momentum through the idea that optimization problems encoding prior domain knowledge can be embedded as parameterized “layers” within neural networks – training such layers requires computing the derivative (gradient) of the optimization problem’s solution during backpropagation [5, 1, 23].
@@
Despite significant research efforts, existing approaches for differentiating QPs fail to fully leverage state-of-the-art solvers. General-purpose methods that support a broad class of optimization problems often include QPs, but they sacrifice efficiency and robustness for generality, underperforming compared to QP-specific methods. On the other hand, recent QP-specific approaches are tightly coupled with particular or proprietary solvers (e.g., for batching small QPs), limiting their flexibility [5, 17, 29].
This tight integration restricts broader applicability. Solving QPs efficiently and reliably often requires advanced solvers like Gurobi [58] and MOSEK [10], developed over years of academic and commercial effort. These solvers handle scales and complexities that non-industrial implementations cannot. However, no single solver is optimal for all problems, making solver selection critical. To our knowledge, no existing approach provides a robust, efficient, and fully flexible framework and implementation for differentiating QPs across solvers.
We introduce dQP, a modular, solver-agnostic framework that transforms any QP solver into a differentiable layer. We strengthen modularity by introducing an explicit characterization of a QP’s primal–dual solution in terms of its active constraint set. Our key insight is that once the active inequalities are known, both the solution and its derivative can be computed from simplified linear systems that share the same matrix. This formulation not only fully decouples optimization from differentiation but also enables differentiation of solvers that output only primal solutions, by completing the corresponding dual variables at negligible computational cost.
We implement dQP as an open-source, minimal-overhead layer on top of qpsolvers [32], enabling seamless integration with over 15 commercial and free solvers, and straightforward extension to additional ones. Our modular approach allows users to select the best solver for their task while providing differentiability. We evaluate dQP on a benchmark of over 2,000 diverse QPs and demonstrate superior performance compared to existing methods, particularly in large-scale structured sparse problems, as demonstrated in Figure 1.
@@
    We design and implement a modular differentiable layer compatible with any QP solver. Our open-source code is available at https://github.com/cwmagoon/dQP.
@@
    We introduce an explicit, simplified characterization of a QP’s solution and gradient in terms of its active constraints, which underpins our modular framework.
Figure 1: Comparison of differentiable QP methods for projection onto the probability simplex, evaluated by (a) total solve and differentiation time and (b) solution accuracy (duality gap). For moderately sized problems, our approach, using the Gurobi QP solver, outperforms existing methods on both metrics. As problem size increases, our method remains efficient, while others become intractable.
@@ 2 Related Works
Differentiable Optimization. OptNet [5, 4] differentiates QPs through their optimality conditions, and focuses on small dense problems for GPU batching. They solve the full Jacobian system efficiently by reusing the factorization employed in their custom interior-point method. However, as noted in [17], this comes at the cost of ill-conditioning due to symmetrization. More recent differentiable QP methods include Alt-Diff and SCQPTH [109, 30, 29] which use first-order ADMM and approximately differentiate the fixed point map, and QPLayer [17] which accommodates infeasibility with extended conservative Jacobians. Similarly to OptNet, several of these are tightly integrated with custom algorithms, often to enable access to internal computations required for differentiation. Alt-Diff is coupled with a custom ADMM method, SCQPTH reimplements OSQP, and QPLayer is built on ProxQP. Several works have noticed the importance of the active constraint set in differentiating constrained optimization problems [8, 57, 89], as well as in the context of quadratic programming [9, 17, 86, 83]. A common observation is that the algebraic system obtained through implicit differentiation can be simplified by removing rows corresponding to inactive constraints. Some works have additionally observed that backpropagation itself can be cast as a distinct equality-constrained QP parameterized by the incoming gradients [8, 86]. However, existing approaches have not used these observations to create a differentiable layer that supports arbitrary black-box QP solvers, missing the opportunity to fully decouple optimization and differentiation. Other classes of optimization problems, such as convex cone programs [2] and mixed-integer programs [89], have also been differentiated. Some frameworks [1, 24, 92, 39, 96, 22, 90, 102] provide differentiable interfaces to broader classes of optimization problems, but their support of QP has significant limitations. CVXPYLayers [1] reformulates the QP into a cone program to use diffcp internally [2] and, as a result, does not support specialized QP solvers, relying exclusively on the cone solvers SCS, ECOS, and Clarabel. The framework Theseus [92] directly handles only unconstrained problems and similarly lacks support for QP-specific solvers. TorchOpt [96] and Optax [39] support unconstrained optimizers used in meta-learning, but require user-defined optimality conditions and user-supplied solvers to handle problems such as QPs. JAXopt [24] includes an implicit differentiation wrapper for CVXPY, which requires symbolic compilation of the QP and does not support sparse matrices; for sparse QPs, JAXopt includes a differentiable re-implementation of OSQP (similar to SCQPTH). Altogether, these drawbacks lead to subpar performance which is reported for some generic methods in previous work on differentiable QPs (e.g., [17]). For completeness, Table 4 in Appendix A summarizes 24 relevant methods, including those discussed above.
@@
Implicit Layers. Optimization layers are a class of implicit layers that use implicit differentiation to compute gradients of solution mappings that lack closed-form expressions [43]. Another class of implicit layer are deep equilibrium models, which are defined by fixed-point mappings and can be interpreted as infinitely deep networks [14, 66, 44, 59, 113, 15]. Similar techniques extend beyond algebraic equations to neural ODEs, where the adjoint state method from parametric PDE control is applied [76, 115, 18, 34]. Implicit differentiation also plays a key role in bi-level programming [36, 72, 56, 3] and meta-learning, where it enables optimization of the outer learning loop [50, 11, 61, 63, 95, 101]. Alternative methods bypass implicit differentiation using approximations – for example, by applying automatic differentiation to iterative algorithms via loop unrolling [19, 20, 81, 105], or by differentiating a single iteration or using Jacobian-free techniques for fixed-point mappings [53, 52, 25].
@@
Sensitivity Analysis and Parametric Programming. There is extensive mathematical theory on the local behavior of optimization problems under perturbations [99, 100], particularly regarding solution sensitivity and stability [47, 49, 48, 73, 26]. While the implicit function theorem is central to this analysis, applying it to optimization problems requires an intermediate step: reformulating the problem through its optimality conditions, which demands several regularity assumptions. Sensitivity theory supports applications such as multi-parametric programming [93], including model predictive control, where problems must often be solved repeatedly for many parameter values, increasing computational costs. To address this, [21] observed that QPs admit closed-form solutions when the active set is known in advance, enabling offline precomputation. The parameter space can be partitioned into regions with fixed active sets [106], where the solution is stable under perturbations. This idea continues to inform modern methods [45, 82, 12].
@@ 3 Approach
We begin by formulating the problem and establishing the basic theory of QP differentiation. Then, we connect this foundation to our central theoretical observation and conclude with our straightforward algorithm derived from it. We note that various subparts of our discussion have been used to develop methods for differentiating QPs, see Section 2. However, no single work has fully developed the explicit theory we present or leveraged it to practically implement a fully modular differentiable QP layer.
@@ 3.1 Problem Setup: Differentiating QPs
This work focuses on computing $\partial_{\theta}z^{*}(\theta)=\frac{\partial}{\partial\theta}z^{*}(\theta)$, the derivative of the optimal point of the QP (1) with respect to the parameters $\theta$. Intuitively, this derivative quantifies the change in the optimal point of the QP in response to a change in its parameters $\theta$. Our goal is to compute $\partial_{\theta}z^{*}(\theta)$ efficiently and independently of the method used to approximate the optimal point $z^{*}(\theta)$.
An important use case, highlighted in recent work on differentiable optimization, involves incorporating a QP of the form (1) as the $\ell$-th layer of a neural network. In this setting, the input to the QP layer, $x_{\ell}$, serves as the parameter vector $\theta$ that defines the QP’s objective and constraints, and the output is the optimal solution $x_{\ell+1}=z^{*}(x_{\ell})$. Generally, training such a network requires backpropagating gradients through each layer, which involves computing the Jacobian $\frac{\partial x_{\ell+1}}{\partial x_{\ell}}$. For a QP layer, this Jacobian corresponds to $\partial_{\theta}z^{*}(\theta)$, the derivative of the optimal point with respect to the problem parameters. This derivative is also crucial in descent-based methods for solving bi-level optimization problems [36].
@@ 3.2 KKT Conditions and Sensitivity Analysis
The first-order Karush–Kuhn–Tucker (KKT) conditions [65, 71, 27, 114] provide a useful algebraic characterization of the optimal points of constrained optimization problems. For the QP (1), the KKT conditions take the form,
@@
To differentiate QPs, we apply the Basic Sensitivity Theorem (Theorem 2.1 in [46]), which underpins differentiation of the KKT conditions with respect to $\theta$. Differentiability at $\theta$ requires the additional assumption of strict complementary slackness, ruling out the degenerate case where both $(Cz^{*}-d)_{j}=0$ and $\mu^{*}_{j}=0$, thus ensuring that the active set remains unchanged under small perturbations of $\theta$. Under this assumption, the primal-dual point $\zeta^{*}(\theta)=(z^{*}(\theta),\lambda^{*}(\theta),\mu^{*}(\theta))$ is differentiable in a neighborhood of $\theta$, optimal for the QP (1), uniquely satisfies the KKT conditions (2), and maintains strict complementary slackness. Crucially, the active set $J(\theta)$ remains fixed in this neighborhood.
@@
Under the assumptions of the Basic Sensitivity Theorem, the linear system (3) is invertible. It degenerates exactly in the presence of weakly active constraints $\mu_{j}^{*}=(Cz^{*}-d)_{j}=0$, for which the QP is non-differentiable (see, e.g., [5]). For any inactive constraint $j\notin J$, the dual variable $\mu_{j}^{*}$ vanishes, and thus the corresponding rows and columns of (3) can be removed, yielding the simplified reduced form,